p+3/4=9/10
I need to be able to show the work and can't seem to get it.

Answer by checkley71(8403)

P+3/4=9/10
P=9/10-3/4 FIND A COMMON DENOMINATOR OF 4 & 10 WHICH IS 20
P=18/20-15/20
P=3/20
PROOF
3/20+3/4=9/10 STILL NEED A COMMON DENOMINATOR 20 WORKS HERE
3/20+15/20=18/20
18/20=18/20